# Inverse Constitutional AI: Compressing Preferences into Principles

### v1

@@ Appendix F Synthetic data generation
In this Appendix, we describe our approach to generate the data used for the experiments in Section 4.1. We use an LLM (gpt-3.5-turbo-0125) to generate the data, given the various prompts below. We include the full data generation pipeline as part of our code, including additional prompt appendices that are added for more diversity. For each of the principles here we generate 10 samples with the corresponding prompts. The unaligned data is simply the aligned dataset with the preference labels flipped.
F.1 Orthogonal data
F.1.1 Example
Listing 30: Orthogonal data example
@@ F.1.1 Example
F.1.2 Prompts for principle "prefer cats over dogs"
Listing 31: Generation prompt 1 for synthetic orthogonal data
@@ F.1.2 Prompts for principle "prefer cats over dogs"
Listing 32: Instruction 1 for synthetic orthogonal data
@@
F.1.3 Prompts for principle "prefer green over blue color"
Listing 33: Generation prompt 2 for synthetic orthogonal data
@@ F.1.3 Prompts for principle "prefer green over blue color"
Listing 34: Instruction 2 for synthetic orthogonal data
@@
F.1.4 Prompts for principle "select lemon over raspberry ice-cream"
Listing 35: Generation prompt 3 for synthetic orthogonal data
@@ F.1.4 Prompts for principle "select lemon over raspberry ice-cream"
Listing 36: Instruction 3 for synthetic orthogonal data
@@
F.2 Aligned data
F.3 Example
Listing 37: Aligned data example
@@ F.3 Example
F.3.1 Prompts for principle "select truthful over factually incorrect answers"
Listing 38: Generation prompt 1a for synthetic aligned data
@@ F.3.1 Prompts for principle "select truthful over factually incorrect answers"
Listing 39: Generation prompt 1b for synthetic aligned data
@@
In the dataset, we replace this original instructions with the following instruction.
Listing 40: Instruction 1 for synthetic aligned data
@@
F.3.2 Prompts for principle "select helpful over useless answers"
Listing 41: Generation prompt 2a for synthetic aligned data
@@ F.3.2 Prompts for principle "select helpful over useless answers"
Listing 42: Generation prompt 2b for synthetic aligned data
@@
In the dataset, we replace this original instructions with the following instruction.
Listing 43: Instruction 2 for synthetic aligned data
@@
F.3.3 Prompts for principle "select polite over impolite answers"
Listing 44: Generation prompt 3a for synthetic aligned data
@@ F.3.3 Prompts for principle "select polite over impolite answers"
Listing 45: Generation prompt 3b for synthetic aligned data
@@
Listing 46: Instruction 3 for synthetic aligned data

### v2

@@ Appendix I Synthetic data generation
In this Appendix, we describe our approach to generate the data used for the experiments in Section 4.1. We use an LLM (gpt-3.5-turbo-0125) to generate the data, given the various prompts below. We include the full data generation pipeline as part of our code, including additional prompt appendices that are added for more diversity. For each of the principles here, we generate 10 samples with the corresponding prompts. The unaligned data is simply the aligned dataset with the preference labels flipped.
I.1 Orthogonal data
I.1.1 Example
Listing 39: Orthogonal data example
@@ I.1.1 Example
I.1.2 Prompts for principle “prefer cats over dogs”
Listing 40: Generation prompt 1 for synthetic orthogonal data
@@ I.1.2 Prompts for principle “prefer cats over dogs”
Listing 41: Instruction 1 for synthetic orthogonal data
@@
I.1.3 Prompts for principle “prefer green over blue color”
Listing 42: Generation prompt 2 for synthetic orthogonal data
@@ I.1.3 Prompts for principle “prefer green over blue color”
Listing 43: Instruction 2 for synthetic orthogonal data
@@
I.1.4 Prompts for principle “select lemon over raspberry ice-cream”
Listing 44: Generation prompt 3 for synthetic orthogonal data
@@ I.1.4 Prompts for principle “select lemon over raspberry ice-cream”
Listing 45: Instruction 3 for synthetic orthogonal data
@@
I.2 Aligned data
I.3 Example
Listing 46: Aligned data example
@@ I.3 Example
I.3.1 Prompts for principle “select truthful over factually incorrect answers”
Listing 47: Generation prompt 1a for synthetic aligned data
@@ I.3.1 Prompts for principle “select truthful over factually incorrect answers”
Listing 48: Generation prompt 1b for synthetic aligned data
@@
In the dataset, we replace these original instructions with the following instruction.
Listing 49: Instruction 1 for synthetic aligned data
@@
I.3.2 Prompts for principle “select helpful over useless answers”
Listing 50: Generation prompt 2a for synthetic aligned data
@@ I.3.2 Prompts for principle “select helpful over useless answers”
Listing 51: Generation prompt 2b for synthetic aligned data
@@
In the dataset, we replace these original instructions with the following instruction.
Listing 52: Instruction 2 for synthetic aligned data
@@
I.3.3 Prompts for principle “select polite over impolite answers”
Listing 53: Generation prompt 3a for synthetic aligned data
@@ I.3.3 Prompts for principle “select polite over impolite answers”
Listing 54: Generation prompt 3b for synthetic aligned data
@@
Listing 55: Instruction 3 for synthetic aligned data